OpenStudy (anonymous):

Convert 3 – 3i to polar form.

OpenStudy (anonymous):

take , r cos theta = 3 r sin theta = -3 now, r^2 (cos^2 theta + sin^2 theta) = ?

OpenStudy (anonymous):

i got 0? @gemini208012

OpenStudy (anonymous):

How did you do it ?

OpenStudy (anonymous):

oh wait 18? I because don't just square the two and add them?

OpenStudy (anonymous):

@gemini208012

OpenStudy (anonymous):

yes so, if r^2 = 18 then, r = root 18

OpenStudy (anonymous):

so my answer would be 18?? @gemini208012

OpenStudy (anonymous):

now put the value of "r" in the first two equations. which is r cos theta = 3 cos theta = ? and r sin theta = -3 sin theta = ?

OpenStudy (anonymous):

I'm sorry I'm confused pre-cal is not my subject! @gemini208012

OpenStudy (anonymous):

cos theta = 3/r cos theta = 3/root 18 and sin theta = -3/r sin theta = -3/root 18 did you get it ?

OpenStudy (anonymous):

eh no I picked 0 I didn't know @gemini208012

OpenStudy (anonymous):

what is it that you did not understand ?

OpenStudy (anonymous):

the whole set up period like I didn't know where to plug in the numbers @gemini208012

OpenStudy (anonymous):

see earlier we had r cos theta = 3 we calculated r^2 = 18 so r = root 18 (by removing square from "r") now to the first equation we had r cos theta = 3 root 18 cos theta = 3 (putting the value of "r") cos theta = 3/root 18 and similarly for sin theta .

OpenStudy (anonymous):

which will be \[\sin \theta = -3/\sqrt{18}\]

OpenStudy (anonymous):

so we have \[\sin \Theta = -3/\sqrt{18}\] \[\cos \Theta = 3/\sqrt{18}\]

OpenStudy (anonymous):

And those cancel out right? @gemini208012

OpenStudy (anonymous):

Yes cancelling will give you \[\sin \theta = -1/\sqrt{2}\] \[\cos \theta = 1/\sqrt{2}\]

OpenStudy (anonymous):

but then those wouldn't cancel either? @gemini208012 that's where I keep getting stuck .

OpenStudy (anonymous):

See you can write \[\sqrt{18} = \sqrt{3 * 3 * 2}\]

OpenStudy (anonymous):

which is \[\sqrt{18} = 3\sqrt{2}\]

OpenStudy (anonymous):

that would be the final answer ? @gemini208012

OpenStudy (anonymous):

no . Now that you have the values . You need to write them in the pi form.

OpenStudy (anonymous):

3pi/4? @gemini208012

OpenStudy (anonymous):

No. See that the sin theta is "minus".

OpenStudy (anonymous):

sin theta is minus in the third quadrant.

OpenStudy (anonymous):

fourth quadrant will also do it.

OpenStudy (anonymous):

lol i'm just confused! @gemini208012

OpenStudy (anonymous):

You can write 7pi/4, but 5pi/4 will also do it. because cos -theta = cos theta

OpenStudy (anonymous):

Now you know the value of r and both sin and cos. Write them in the form of r(cos theta + i sin theta)

OpenStudy (anonymous):

I give up lol @gemini208012

OpenStudy (anonymous):

it will be \[\sqrt{18} (\cos 7\Pi/4 + i \sin 7\Pi/4)\]

OpenStudy (anonymous):

Did you get it ?

OpenStudy (anonymous):

Okkkkkkay I kinda see what your saying now @gemini208012

OpenStudy (anonymous):

lol it is not that hard. Just practice. I was also confused like you the first time I solved them.

OpenStudy (anonymous):

You helped me out though thank you so much! @gemini208012

OpenStudy (anonymous):

You are welcome. @Tashaydavis63